Corporation purchased a printing machine three (3) years ago and is considering replacing it with a new one which is faster and easier to operate. The old machine has been depreciated over 3 years using straight-line depreciation. Its original installation cost was $15,000. The old machine has been in use for 2 years, and it can be traded in for $3,500. The new machine will be purchased $24,000 and it will also be depreciated over 3 years using the straight-line method. It is not expected to have a residual value. Networking capital will decrease because supply levels can be reduced by $1,500. Revenues will increase by $5,000 every year, it will result in labor savings of $3,000 per year due to its greater speed. Reducing training expenses is expected to save an additional $2,500 per year. The firm is in the 20% tax bracket.
Answer #1

Calculation:

=> EBITDA = 5000 + 3000 + 2500 = 10500

Operating cash flow formula =  EBITDA ( 1 - t ) + Depreciation x t + changes in net working capital

=> $ 10,500 x 0.8 + $ 8,000 x 0.2 - 1500 = $ 10,000.

Terminal non-operating cash flow = Increase in Working Capital = - $ 1,500.